OpenStudy (anonymous):

y= -1/2x-1 what is the equation od the line that is perpendicular to the first line and passes through the point ( 2, -5)?

OpenStudy (anonymous):

For two lines to be perpendicular m1*m2=-1

OpenStudy (anonymous):

m=2 -5=4+b b=-9 y=2x-9